# LLM-enhanced Scene Graph Learning for Household Rearrangement

### v1

@@ 3.2. Misplaced Object Detection
Since our task is to “tide the house”, we are looking for carriables that currently have an unreasonable placement and need to be rearranged to complete the task. To achieve this, we introduced a LLM-based placement scorer to rate the suitability of a placement based on the affordance and description of carriable-receptacle pair. With this placement scorer, we can determine if a carriable object is appropriately placed currently and find all the misplaced carriables.
Specifically, Due to the token limitation of the LLM, we cannot input all the carriable objects at once for scoring, so we score carriables one by one. For each carriable object, we collect the description of the carriable object, the context-induced affordance of the receptacle it is currently located, and our task instruction, which is ”tide the house.” With the aboved collected information as well as a well-designed prompt, LLM will analyze the current placement of the given carriable object to determine the suitability of the placement, and assign a score to this relevance.
Note that scoring objects individually can lead to variability in standards due to the LLM’s lack of criteria context. To mitigate this, we set a fixed standard in the prompt for the LLM to refer to when rating: 100 points: the placement perfectly meets the task requirements; 0 points: the placement contradicts the task requirements and might negatively impact the task, and thus need to be rearranged; 50 points: It is difficult to judge whether rearranging this object is related to the task. Figure 5 shows an example of the scoring process. We set 50 points as the threshold and all those carriable objects with score no more than 50 are considered to be misplaced and should be rearranged. These carriables are then reorganized in accordance with their respective scores, arranged in ascending order, to facilitate the planning of subsequent rearrangements.
@@ 3.3. Object Rearrangement Planning
For each misplaced object, our goal is to find the most suitable receptacle in the scene as the placement target with the context-induced affordance information stored in the AEG. To allow LLM to find the most suitable placement, it is necessary to consider all the relevant receptacles for comparison. However, directly inputting the affordances of all possible receptacles in the AEG to LLM will not only flood it with excessive irrelevant information, impairing its reasoning ability, but also makes it lose focus, increasing the likelihood of hallucinations.
Inspired by Retrieval Augmented Generation (RAG) (Gao et al., 2023a), which retrieves the task-related text chunks from a given database to augment LLM’s task-solving ability and prevent hallucination, we construct a database of receptacle affordance and retrieve the top-k task-related receptacles from this database as suitable candidates and send them to LLM to select the best target, as shown in Figure 6. In this way, the task-irrelevant receptacles as well as the distracting information will be filtered during the retrieving process, and with the rest top-k candidates as prompt, LLM can make accurate placement decision and generate high-performance rearrangement planning.
@@
Different from the classic RAG, which splits a long text into multiple chunks to create the database and retrieves the most relevant chunks based on semantic similarity to the query, our dataset consists of descriptions of all receptacle nodes in the AEG, and the top-k most relevant receptacles are retrieved based on task relevance to the query.
To achieve this, we adopt our LLM scorer from misplacement detector with a minor change of prompt and use it to rate the appropriateness of placement based on task instruction. The inputs are same as the misplacement detector, consisting of the task instruction “tide the house”, the carriable object description, as well as the receptacle’s context-induced affordance. However, the scoring standards in the prompt are adjusted accordingly as follow: “ 100 points: The LLM believes this receptacle is the most suitable place in the entire house to place the carriable in order to complete the task; 0 points: The LLM believes that, from the task’s perspective, placing the given carriable in this receptacle should not be considered under any circumstances; 50 points: The LLM believes that placing the object in this receptacle could only be reasonable under certain special conditions. ” With all receptacle rated, we can choose the top k receptacles with the highest scores as candidates and use their context-induced affordance as retrievals for the placement decision generation.
@@
We introduce a LLM-based placement decision generator to select the best receptacle among the proposed candidates and generate rearrangement plan. With the top-k rated receptacles, we feed their context-induced affordance as well as the query (description of carriable to rearrange together with the task) into the placement decision generator. The LLM then selects the best placement target, outputting the placement decision as following format: “ Best receptacle: the name of the most suitable receptacle among the candidates; Analysis: the reason why it thinks this receptacle is the best one compared to others.
With the generated placement target, we can pick up all the misplaced carriable objects and transfer them to their target placement location, completing the rearrangement process.
@@ 4. Experiment and Results
4.1. Benchmark
Datasets.
Two benchmarks were used for testing, one proposed by Tidybot (Wu et al., 2023) and another dataset annotated by us based on the Habitat Synthetic Scenes Dataset (HSSD 200) (Khanna et al., 2023). It is worth noting that the scenes in the Tidybot dataset are all single-room, with generally no more than 5 objects per scene and relatively sparse context information. Therefore, to fully validate the performance of methods in scenarios closer to real-world settings, we established a context-oriented dataset consisting of 8 scenes based on existing complex scene datasets, totaling 36 rooms, 105 areas, and 188 receptacle objects. We randomly place carriable objects into the benchmark scenarios to generate more than 4000 messy scenes for misplacement detection evaluation.
@@ 4.1. Benchmark
For the household rearrangement task, we evaluated our method mainly in two aspects: misplacement detection and carriable object rearrangement planning. When the ground-truth object placements are available, we measure the detection performance based on accuracy, recall, precision, and F1 score following most existing works (Padilla et al., 2020). To evaluate the performance of carriable object rearrangement planning, we take the best 8 placements for each carriable object to calculate the Normalized Discounted Cumulative Gain (NDCG) based on the labeled ground-truth receptacles. More details about the benchmark and evaluation metrics can be found in the supplementary material.
@@ 4.2. Quantitative Comparisons
Note that Tidybot fails when testing on our context-oriented benchmark due to its limited performance in large-scale scenes and the setting where no certain demonstration placements are provided. Thus, we compare our method with a state-of-the-art alternative Housekeep (Kant et al., 2022), a learning-based rearranging method that learns the knowledge priors from the semantic context of the existing scenes. As shown in the comparison results in Table 2, our approach significantly outperforms Housekeep in both placement rearrangement and misplacement detection tasks. It demonstrates that the proposed AEG can significantly better enhance the agent’s understanding of scenes by leveraging additional context such as spatial, geometric, and visual cues besides semantic relationships.
@@ 4.3. Ablation Studies
Necessity of LLM-based scorer.
Compared to asking the LLM to predict the final receptacle directly based on AEG, we find that adopting an LLM-based scorer to retrieve candidates for an RAG query can provide more reasonable results. To make a fair evaluation, we compare our LLM-based scorer strategy with a similarity-based candidate retrieval method (Gao et al., 2023b) that is widely used in NLP applications. We also added a random selection strategy that serves as a baseline. As presented in the right of Figure 7, our LLM-based scorer can provide significantly better performance than other alternatives. We can effectively filter out a large number of irrelevant options with the LLM-based scoring, which greatly reduces the difficulty for the LLM in making the final decision.
@@ 4.4. Qualitative Results
To better illustrate the pipeline of our method, we show some examples of context-induced affordance enhancement in Figure 8. Note how the agent derives more accurate functional descriptions of target objects than the original general semantic tags. Besides, we showcase several examples of carriable object placement planning based on AEG in Figure 9 with real-world scenarios and Figure 10 with synthetic environments. We can see that the agent can plan object placements that better fit the scene context based on AEG. More results can be found in supplementary material.

### v2

@@ 3.2. Misplaced Object Detection
Since our task is to “tide the house”, we are looking for carriables that currently have an unreasonable placement and need to be rearranged to complete the task. To achieve this, we introduced an LLM-based placement scorer to rate the suitability of a placement based on the affordance and description of carriable-receptacle pair. With this placement scorer, we can determine if a carriable object is appropriately placed currently and find all the misplaced carriables.
Specifically, Due to the token limitation of the LLM, we cannot input all the carriable objects at once for scoring, so we score carriables one by one. For each carriable object, we collect the description of the carriable object, the context-induced affordance of the receptacle where it is currently located, and our task instruction, which is ”tide the house.” With the collected information above as well as a well-designed prompt, LLM will analyze the current placement of the given carriable object to determine the suitability of the placement and assign a $0\textasciitilde 100$ score to this relevance. Figure 5 shows an example of the scoring process. We set 50 points as the threshold and all those carriable objects with scores no more than 50 are considered to be misplaced and should be rearranged. These carriables are then reorganized in accordance with their respective scores, arranged in ascending order, to facilitate the planning of subsequent rearrangements.
@@ 3.3. Object Rearrangement Planning
For each misplaced object, our goal is to find the most suitable receptacle in the scene as the placement target with the context-induced affordance information stored in the AEG. To allow LLM to find the most suitable placement, it is necessary to consider all the relevant receptacles for comparison. However, directly inputting the affordances of all possible receptacles in the AEG to LLM will not only flood it with excessive irrelevant information, impairing its reasoning ability but also make it lose focus, increasing the likelihood of hallucinations. Inspired by Retrieval Augmented Generation (RAG) (Gao et al., 2023a), which retrieves the task-related text chunks from a given database to augment LLM’s task-solving ability and prevent hallucination, we construct a database of receptacle affordance and retrieve the top-k task-related receptacles from this database as suitable candidates and send them to LLM to select the best target, as shown in Figure 6. In this way, the task-irrelevant receptacles will be filtered during the retrieving process, and with the rest top-k candidates as prompt, LLM can make accurate placement decisions and generate high-performance rearrangement planning.
@@
Different from the classic RAG, which splits a long text into multiple chunks to create the database and retrieves the most relevant chunks based on semantic similarity to the query, our dataset consists of descriptions of all receptacle nodes in the AEG and the top-k most relevant receptacles are retrieved based on task relevance to the query. To achieve this, we adopt our LLM scorer from the misplacement detector with a minor change of prompt and use it to rate the appropriateness of placement based on task instruction. With all placements of receptacle rated, we can choose the top k receptacles with the highest scores as candidates and use their context-induced affordance as retrievals for the placement decision generation.
@@
We introduce an LLM-based placement decision generator to select the best receptacle among the proposed candidates and generate a rearrangement plan. With the top-k rated receptacles, we feed their context-induced affordance as well as the query (description of carriable to rearrange together with the task) into the placement decision generator. The LLM then selects the best placement target, outputting the placement decision with a paragraph of analysis. With the generated placement target, we can pick up all the misplaced carriable objects and put them in their target locations, completing the rearrangement process.
@@ 4. Experiment and Results
4.1. Benchmark
Datasets.
Two benchmarks were used for testing, one proposed by Tidybot (Wu et al., 2023) and another dataset annotated by us based on the Habitat Synthetic Scenes Dataset (HSSD 200) (Khanna et al., 2023). The TidyBot [Wu et al. 2023] benchmark is centered around learning individual preferences from human-provided examples, which results in smaller scenes with each object having a single ground truth (GT) placement. In contrast, our focus is on inferring commonsense object arrangements in large, house-level scenes, which offer richer contextual information and more varied placement options. To fully assess performance in scenarios that better resemble real-world conditions, we developed a context-driven benchmark comprising 8 scenes from Habitat Synthetic Scenes Dataset (HSSD 200) [Khanna et al. 2023], including 36 rooms, 105 areas, and 188 receptacle objects. We randomly placed carriable objects in these scenes, creating over 4000 messy scenes to evaluate misplacement detection. To determine correct object placement, we enlisted skilled individuals to control a robot, navigating each scene to identify the most appropriate receptacle for each new object, with a minimum of one and a maximum of five options. As all scene information was readily accessible, the provided annotations accurately reflect human commonsense about object placement. After gathering all annotations, we calculated the majority vote for each receptacle per object in each scene, generating a ranked list of receptacle choices.
@@ 4.1. Benchmark
For the household rearrangement task, we evaluated our method mainly in two aspects: misplacement detection and carriable object rearrangement planning. When the ground-truth object placements are available, we measure the detection performance based on accuracy, recall, precision, and F1 score following most existing works (Padilla et al., 2020). In our benchmark tests, each object may have multiple placement positions, with a ranking that reflects human common sense. To more comprehensively compare the performance of different methods in carriable object rearrangement planning, we also require these methods to output a ranking of recommended placement positions. Normalized Documented Cumulative Gain, which is regarded as a measure of how close a ranking is to the ground truth ranking, can assess whether the methods can select appropriate placement positions based on human common sense in the rearrangement planning. In practice, we take the best 8 placements for each carriable object to calculate the NDCG.
@@ 4.2. Quantitative Comparisons
Note that Tidybot fails when testing on our context-oriented benchmark due to its limited performance in large-scale scenes and the setting where no certain demonstration placements are provided. Thus, we compare our method with alternative Housekeep (Kant et al., 2022), a state-of-the-art learning-based rearranging method that extracts priors from the joint probability distribution among rooms, receptacles, and objects provided by their human-annotated dataset. Unlike our method, the approach does not consider the functionality implied by the spatial relationships of surrounding objects and the objects’ geometries. As shown in the comparison results in Table 2, our approach significantly outperforms Housekeep in both placement rearrangement and misplacement detection tasks. It demonstrates that the proposed AEG can significantly better enhance the agent’s understanding of scenes by leveraging additional context such as spatial, geometric, and visual cues besides semantic relationships.
@@ 4.3. Ablation Studies
Necessity of LLM score-based retriever.
Compared to asking the LLM to predict the final receptacle directly based on AEG, we find that adopting an LLM scorer to retrieve candidates for a RAG query can provide more reasonable results. To make a fair evaluation, we compare our LLM score-based retrieval strategy with a similarity-based candidate retrieval method (Gao et al., 2023b) that is widely used in NLP applications. We also added a random selection strategy that serves as a baseline. As presented in the right of Figure 7, our LLM score-based retriever can provide significantly better performance than other alternatives. We can effectively filter out a large number of irrelevant options with the LLM-based scoring, which greatly reduces the difficulty for the LLM in making the decision.
We also implemented ablations about key frame selection, number of retrievals, and score prompting in the supplementary material.
@@ 4.4. Qualitative Results
To better illustrate the pipeline of our method, we show some examples of context-induced affordance enhancement in Figure 8. Note how the agent derives more accurate functional descriptions of target objects than the original general semantic tags. Besides, we showcase several examples of carriable object placement planning based on AEG in Figure 9 with real-world scenarios and Figure 10 with synthetic environments. We can see that the agent can plan object placements that better fit the scene context based on AEG.